The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If 485 out of the next 500 are processed on the same day, would this prove that they are exceeding their goal, using alpha.025?
Answer by stanbon(75887)

The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If 485 out of the next 500 are processed on the same day, would this prove that they are exceeding their goal, using alpha.025?
--------------------------------------
Ho: p = 0.95
Ha: p > 0.95 Claim
--------------------
Critical Value for one-tail z-test with alpha = 0.25 = 1.96
-------------------
Test statistic: z(485/500) = (0.97-0.95)/sqrt[0.95*0.05/500] = 2.05195...
---------------------
p-value = P(2.05195 < z < 10) = 0.02
----------------------
Conclusion: Since p-value is less than alpha, Reject Ho:
The test results support the Claim.
==================
Cheers,
Stan H.